The tensile strength of a metal part is normally distributed with mean 35 pounds and standard deviation 5 pounds. Suppose 40,000 parts are produced and specifications on the part have been established as 35.0 ± 4.2 pounds.

Find the percentage of parts that will fail to meet specification

Find the number of parts that will fail to meet specifications

Find the tensile strength at which 10% of the parts exceed the upper specification limit
Answer #1

onscoes! Let Mean - dl= 35 Standard deviation=6 - 5 UCL = 35.0+4,2 LEL=35.0-41.2 - 39.2 - 30.8 a) Picolul fail to meet specifb) Let No.of parts that coill fail to meet specifications *** : 40,000 loucoq). 40,000 (0.4009). 16036 colere. P(x>0) = { 17